# Semantic Inference-Based Deep Learning and Modeling for Earth Observation: Cognitive Semantic Augmentation Satellite Networks

### v1

@@ V-A4 On-board processing efficiency
Transformer-based models, such as ViT, SwinTransformer, and EfficientNet variants, are far more computationally intensive. ViT-large (Google) leads the complexity spectrum with 303 million parameters, 162.20 GFLOPs, and significant memory requirements (1.24 GB), translating into much longer training (9,271 s) and inference (2 s) times. SwinTransformer and EfficientNet-V2 models offer a compromise between complexity and efficiency. SwinTransformer, for instance, has 85.7 million parameters, 21.57 GFLOPs, and takes around 538 seconds for training, making it efficient for high-performance tasks despite its complexity. The authors highlights that transformer models like ViT-huge (Google) and ViT-large are the most resource-intensive, while smaller models like SmallViT and EfficientNet-V2 maintain a good balance between model size, computational complexity, and time efficiency. Ultimately, this comparison shows a trade-off between performance and computational demand, with traditional CNN models being more efficient for simpler tasks, and transformer models excelling at complex, large-scale tasks but requiring significantly more computational resources.
V-B Numerical Results
V-B1 Channel model
Let $X$ denote the transmitted signal and $Y$ denote the received signal. The general form of the channel equation is:$Y=H\cdot X+N$, where $H$ is the channel gain and $N$ is the noise term. Below are the specific formulations for different types of channels.
For an AWGN channel, there is no fading, so the channel gain $H=1$: $Y=X+N_{\text{AWGN}}$, where $N_{\text{AWGN}}\sim\mathcal{N}(0,\sigma^{2})$ is the Gaussian noise with variance $\sigma^{2}$. In a Rician channel, the channel gains $H$ follows a Rician distribution: $Y=H_{\text{Rician}}\cdot X+N$, where $H_{\text{Rician}}$ is a Rician-distributed random variable characterized by the Rician $K_{r}$factor, and $N\sim\mathcal{N}(0,\sigma^{2})$ is the noise. For a Rayleigh channel, the channel gain $H$ follows a Rayleigh distribution: $Y=H_{\text{Rayleigh}}\cdot X+N$, where $H_{\text{Rayleigh}}$ is a Rayleigh-distributed random variable, and $N\sim\mathcal{N}(0,\sigma^{2})$ is the noise.
In the LEO Rician channel, the received signal incorporates the Rician fading characteristics along with effects like Doppler shift and satellite-specific parameters: $Y=H_{\text{LEO-Rician}}\cdot X+N_{\text{AWGN}}$, where $H_{\text{LEO-Rician}}$ is a Rician-distributed fading gain in the LEO environment and includes noise contributions such as scintillation and atmospheric loss according to Table II. For a LEO Rayleigh channel, the fading gain follows a Rayleigh distribution with LEO-specific noise characteristics: $Y=H_{\text{LEO-Rayleigh}}\cdot X+N_{\text{AWGN}}$, where $H_{\text{LEO-Rayleigh}}$ is a Rayleigh-distributed fading gain.
To simulate propagation and attenuation, the channel model incorporates both small- and large-scale fading; the path loss components are the large-scale fading between the satellite and UT [74]: $P_{L_{\text{tot}}}=P_{L_{b}}+P_{L_{g}}+P_{L_{s}}$, in which the attenuation resulting from atmospheric gases is denoted by $P_{L_{g}}$, the attenuation resulting from ionospheric or tropospheric scintillation by $P_{L_{s}}$, and the basic path loss is represented by $P_{L_{b}}$. Furthermore, the ISL of free space may be understood as $P_{L_{\text{ISL}}}=P_{L_{ib}}$, in which $P_{L_{ib}}$ denotes Sat1 and Sat2’s fundamental path loss. These route loss components are all expressed in decibels (dB). More specifically, the free space propagation and shadow fading of the signal are taken into account using the fundamental path loss model $P_{L_{b}}$. For a distance $d_{k,i}$ also known as slant range, the free space path loss ($P_{L_{\text{FS}}}$) in dB range in meters and frequency $f_{c}$ in GHz. Hence, the slant distance $d_{k,i}$ can be expressed as [75]:
$d=\sqrt{R_{E}^{2}\sin^{2}\theta+r_{m}^{2}+2R_{E}r_{m}-2R_{E}r_{m}\sin\theta},$ (1)
where $R_{E}=6378\,\text{km}$ is the radius of the Earth, $r_{m}$ and $\theta$ are the satellite altitude and its elevation angle, respectively.
Thereby, the free space path loss ($P_{L_{\text{FS}}}$) can be calculated as: $P_{L_{\text{FS}}}(d,f_{c})=32.45+20\log_{10}(f_{c})+20\log_{10}(d)$, Shadow fading (SF) is modeled by a log-normal distribution as $\mathcal{N}(0,\sigma^{2}_{SF})$ with zero-mean and $\sigma_{SF}$ standard deviation. The values of $\sigma^{2}_{SF}$ can be extracted from the 3GPP Release-15. Then, the path loss with and without shadow fading in dB units is modeled as $P_{L_{b}}=P_{L_{\text{FS}}}(d,f_{c})+SF$ and $P_{L_{ib}}=P_{L_{\text{FS}}}(d_{ib},f_{c})$, where $d_{ib}$ represent the distance of Sat1 and Sat2.
We consider the Rician model to express the channel between LEO satellite $s1$ and the UT $z$ and ISL between LEO satellite $s1$ and $s2$. The channel from the LEO satellite $s1$ to the UT $z$ and ISL are given as
$f_{s1,z}=\sqrt{\frac{R_{z}\zeta_{s1,z}}{R_{z}+1}}\bar{f}_{s1,z}+\sqrt{\frac{\zeta_{s1,z}}{R_{z}+1}}\tilde{f}_{s1,z},$ (2)
$f_{s1,s2}=\sqrt{\frac{R_{s2}\zeta_{s1,s2}}{R_{s2}+1}}\bar{f}_{s1,s2},$ (3)
where $\bar{f}_{s1,z}$, $\bar{f}_{s1,s2}$ and $\tilde{f}_{s1,z}$ are the line-of-sight (LoS) and non-line-of-sight (NLoS) paths components, respectively, $R_{z}$ is the Rician factor, and the ISL does not have NLoS component. After accounting for the LEO satellite antenna gain $G_{T}$, the large-scale fading effects of the LEO satellite $s1$ to the UT $z$ and ISL may be modeled as follows: $\zeta_{s1,z}(\text{dB})=PL_{\text{tot}}-G_{T}$ and $\zeta_{s1,s2}(\text{dB})=PL_{\text{ISL}}-G_{T}$. Furthermore, the complex Gaussian distribution with zero mean and unit variance is followed by the independent and identically distributed (i.i.d.) entries of $\tilde{f}_{s1,z}$ data.
The complex baseband downlink space domain channel at a time instance $t$ and frequency $f$ can be expressed as follows using a ray-tracing-based channel modeling method:
$H_{s1,z}(t,f)=f_{s1,z}\exp\left(2\pi\left[tv_{z}-f\tau_{z}\right]\right),$ (4)
$H_{s1,s2}(t,f)=f_{s1,s2}\exp\left(2\pi\left[tv_{s2}-f\tau_{s2}\right]\right),$ (5)
where $v_{z}$, $v_{s2}$ are the Doppler shift, and $\tau_{z}$, $\tau_{s2}$ are the propagation delay. Table II summarises the simulation parameters.
TABLE II: Simulation Parameters
Parameter Value
Carrier frequency 28 GHz
Satellite antenna gain 35 dBi
User antenna gain 37 dBi
Scintillation loss 0.5 dB
Atmospheric loss 0.3 dB
Rician factor 2.8
V-B2 Physical layer transmission
Fig. 6 and Fig. 7 present Top-1 accuracy using DT-JSCC under different channel conditions for EuroSAT data. The top two graphs depict performance across different PSNR (Peak Signal-to-Noise Ratio) levels for the 16PSK modulation scheme. We show accuracy for the AWGN (Additive White Gaussian Noise) and Rician channels, with different K-factor values (32, 64, 128), where a higher K-factor generally corresponds to better channel quality. Furthermore, we compare accuracy across the Rician and Rayleigh channels for the same K values, showing higher accuracy in the Rician channel. The lower two graphs show similar comparisons but for the LEO environment with 16APSK modulation. We examine the AWGN and Rician channels, while the right-hand side evaluates Rician and Rayleigh channels under different K-factors. Across all scenarios, the graphs indicate that increasing PSNR improves accuracy, with Rician channels generally performing better than Rayleigh, and higher K-factors leading to more stable performance, especially in LEO environments.
@@ V-B2 Physical layer transmission
Figure 7: Top1 accuracy using DT-JSCC based on 16PSK/16APSK LEO Rician and LEO Rayleigh channel.
In comparing the confusion matrices of Fig.8, we can observe some key performance differences over the AWGN channel where PSNR=14dB. For k=128, the classification accuracy across most categories is notably high, with certain categories like ”Forest” showing a strong prediction rate of 99.21%, while there are minor misclassifications in categories like ”PermanentCrop” and ”Pasture” with accuracies of 93.71% and 94.59% respectively. Meanwhile, the case of K=32 also performs well but demonstrates slightly different classification behaviors, such as a stronger prediction for ”River” at 98.05% but more substantial confusion in categories like ”Forest” and ”Residential,” where there are larger misclassification rates with accuracies of 95.44% and 90.88% compared to the K=128 performance. The identical trend of K=32 and K=128 can be observed over the 16APSK Rician channel. However, Fig.9 demonstrates better accuracy and classification than 16PSK modulation in almost all categories, while there are minor misclassifications in categories like ”HerbaceousVegetation” and ”PermanentCrop” with accuracies of 98% and 98.34% respectively. It is noteworthy that 16APSK modulation validated where PSNR at 12dB is better than 16PSK at 14 dB.
@@
V-B3 CSA satellite network
Table III shows CSA DT-JSCC system with high accuracy and evidently with most classes, such as Annual Crop (94.51%), Herbaceous Vegetation (91.43%), and Sea/Lake (98.85%), showing high Top1 accuracy, indicating correct classifications. Some misclassifications can be improved, such as between Industrial and Residential areas. The non-CSA shows slightly reduced accuracy, with some increased misclassifications, particularly in Herbaceous Vegetation (89.64%) and Sea/Lake (96.63%). Overall, the CSA-enhanced system demonstrates better accuracy and fewer misclassifications compared to the non-CSA system.
TABLE III: Top1 accuracy comparison of CSA and non-CSA DT-JSCC K=32 systems while PSNR=12dB and 16APSK over LEO Rician channel.
Class Accuracy (%) CSA Accuracy (%) non-CSA
AnnualCrop 94.50 88.38
Forest 98.41 98.84
HerbaceousVegetation 91.63 89.64
Highway 92.30 77.27
Industrial 95.95 90.95
Pasture 94.59 90.53
PermanentCrop 90.03 83.55
Residential 98.02 95.95
River 92.33 72.63
SeaLake 98.95 98.33
As illustrated in Fig. 10, the SA technique struggles significantly under the conditions of a low PSNR value due to the inherent noise and interference present in LEO Rician channels. This adverse environment leads to diminished signal quality, which adversely impacts the performance of the SA method. In contrast, a notable enhancement in Top1 accuracy is evident with both K=32 and K=128 DT-JSCC transmissions. This improvement can be attributed to the end-to-end CSA architecture, which leverages advanced techniques to optimize signal processing, resulting in a more robust transmission. By intelligently aggregating information, the system can better maintain signal integrity, thereby improving overall accuracy in challenging channel conditions.
Figure 10: Top1 accuracy of CSA satellite network using DT-JSCC over 16APSK LEO Rician channel.
The results highlight the importance of employing sophisticated transmission strategies, particularly in environments characterized by high levels of noise and interference. The substantial gains in accuracy achieved with the DT-JSCC approach underscore the efficacy of incorporating channel-aware mechanisms in the design of communication systems, suggesting that further exploration and development in this area could yield even more significant benefits in future applications.
@@ VI Conclusion
In this paper, we explored the critical role of semantic inference and deep learning in enhancing EO systems, focusing on the integration of data-driven modeling approaches. With the rise of big EO data and the increasing complexity of satellite imagery, conventional methods struggle to meet the demands of modern applications such as environmental monitoring, disaster management, and resource allocation. By utilizing advanced deep learning techniques, we have illustrated the potential to overcome key challenges, including the scarcity of labeled data, the variability in data sources, and the intricacies of multi-dimensional datasets. The incorporation of semantic inference enables systems to extract meaningful insights from vast and complex datasets, paving the way for more accurate decision-making in both civilian and critical operations.
This work presents an advanced framework for semantic communication networks, with a focus on the development and optimization of Cognitive Semantic EO systems. By incorporating semantic-aware methodologies like JSCC and DT-JSCC, the proposed system enhances communication efficiency, particularly in satellite networks. The integration of semantic data augmentation, inter-satellite links, and cognitive processing capabilities ensures that only the most relevant, task-specific data is transmitted, significantly reducing communication overhead and improving system performance. Furthermore, by leveraging semantic inference and cognitive augmentation techniques, the system facilitates better image analysis, change detection, and decision-making in real-time applications. The proposed end-to-end model demonstrates how integrating semantic cognition into satellite systems can drastically improve data transmission, interpretation, and overall system efficiency, thereby addressing the demands of 6G networks and beyond. This research lays the groundwork for future advancements in semantic communication, with promising applications in a wide range of fields, including remote sensing, autonomous systems, and smart communication networks.
Acknowledgment
This work was funded by the Luxembourg National Research Fund (FNR), with the granted SENTRY project corresponding to grant reference C23/IS/18073708/SENTRY.

### v3

@@ V-A Revisit on the effectiveness of on-board EO deep learning
A detailed comparison of several deep learning models is presented in [71] focusing on metrics like input size, total parameters, estimated model size, FLOPs, and the time taken for training and inference. Simpler architectures, such as CNN and ResNet-50, exhibit low computational complexity. For instance, CNN has the smallest total parameters (66,330) and FLOPs (0.93 MFLOPs), making it highly efficient for training (233 seconds) and inference (7 seconds). ResNet-50, while more complex, still maintains moderate efficiency, with 0.2 million parameters, a model size of 10 MB, and 117.88 MFLOPs, resulting in faster inference times (6.7 s) compared to more complex models. Transformer-based models, such as ViT, SwinTransformer, and EfficientNet variants, are far more computationally intensive. ViT-large (Google) leads the complexity spectrum with 303 million parameters, 59.7 GFLOPs, and significant memory requirements (1.64 GB), translating into much longer training (9,271 s) and inference (60 s) times. SwinTransformer and EfficientNet-V2 models offer a compromise between complexity and efficiency. SwinTransformer, for instance, has 86.7 million parameters, 21.47 GFLOPs, and takes around 538 seconds for inference, making it efficient for high-performance tasks despite its complexity. The authors presented in [71] highlight that transformer models like ViT-huge (Google) and ViT-large are the most resource-intensive, while smaller models like SmallViT and EfficientNet-V2 maintain a good balance between model size, computational complexity, and time efficiency.
Ultimately, this comparison shows a trade-off between performance and computational demand, with traditional CNN models being more efficient for simpler tasks, and transformer models excelling at complex, large-scale tasks but requiring significantly more computational resources.
V-B On-board processing hardware for EO semantic communication
Telecommunication satellites are facing the need to meet the growing demand for faster data rates, smaller antenna sizes, and reduced power consumption in mobile devices. This necessitates the development of a streamlined bridge that is specifically designed for the processing technologies used in the application [72]. On satellites designed for specific purposes or with a limited lifespan, there is some utilization of on-board processing.
V-B1 SRAM-aided programmable on-board processing
The Cibola Flight Experiment (CFE) presented in [73] assessed the feasibility of using SRAM-based Field-Programmable Gate Arrays (FPGAs) for in-flight data processing in spacecraft. Introduced in 2007, it transported three adaptable computer modules, each equipped with Xilinx Virtex XQVR1000 FPGAs for the purpose of data processing. The objective of the experiment was to analyze the impact of radiation on these Field-Programmable Gate Arrays (FPGAs) and assess methods to minimize or counteract these effects. The success of the CFE indicates that SRAM-based FPGAs have promise for future uses in spaceflight. In addition, SRAM-based Field-Programmable Gate Arrays (FPGAs) have shown their worth in performing crucial functions in the field of space exploration [73]. Xilinx XQR4062XL FPGAs were responsible for managing the descent and landing of the Mars rovers Spirit and Opportunity. These missions aimed to overcome the susceptibility of SRAM-based FPGAs to space radiation by using fault mitigation mechanisms. This included implementing Triple Modular Redundancy throughout the design phase and including redundant Field-Programmable Gate Arrays as components. The FPGAs have the capability to be fully reprogrammed in order to effectively eliminate any faults. Although redundancy provides resilience, it also incurs additional weight, power use, and total expenditure.
The Fraunhofer aboard-Board Processor (FOBP) presented in [74] takes a distinct approach by using radiation-hardened Virtex-5QV FPGAs to provide adaptable communication processing aboard satellites. The main benefit of FOBP is its capacity to be reconfigured while in orbit. By reprogramming the FPGAs from ground stations, the processor can adapt to new communication protocols and changes in the environment, going beyond the limits of traditional systems that can not be changed. This opens up the possibility of using non-radiation-protected FPGAs such as Kintex UltraScale or Zynq-UltraScale+ in Low Earth Orbit (LEO) space applications [75]. A method improves onboard reconfiguration of SRAM-based FPGAs in LEO satellite networks [76]. This approach automates the process, breaking down data for transfer, verifying temporary storage, and ensuring successful flash memory programming. Built-in error checks throughout this closed-loop system enhance reliability, allowing for flexible and dependable maintenance of spacecraft FPGAs.
V-B2 Multiprocessor systems-on-chips on-board processing
Multiprocessor Systems-on-Chips (MPSoCs), which integrate CPUs, FPGAs, and other components into a single chip, provide distinct adaptability for space applications. The reason for this is that they integrate the flexibility of software with the programmability of FPGAs. This enables the construction of radiation-resistant systems by using a variety of these structures. An example of this methodology is the CHREC Space Computer (CSP). The CSP presented in [77] employs a Xilinx Zynq platform and a blend of software and hardware hardening methods to attain exceptional dependability, as confirmed by space missions. Like the CSP, the APEX-SoC system similarly uses a single Zynq device for gathering and analyzing data [78].
The potential of reconfigurable MPSoCs is applicable even to space agencies, such as the European Space Agency’s OPS-SAT program [79]. This program places a high priority on using commercially available technologies (COTS) that are already in existence to develop a satellite that is adaptable, allowing for the interchangeability of applications even after it has been launched. The OPS-SAT payload utilizes the ALTERA Cyclone V processor as its core, whereas the NINANO platform [80] employs a Xilinx Zynq-7030 for on-board processing in scientific missions such as EYE-SAT which involves imaging of zodiacal light [81]. The authors of [82] present a specialized on-board CPU architecture that is adaptable and intended for crucial space applications. The primary characteristic is a multi-accelerator hardware system that has the ability to be reprogrammed in real-time while in operation. This enables adaptable optimization of performance, dependability, and energy use according to the requirements of the task. The architecture seamlessly integrates with a run-time reconfigurable operating system, ensuring adherence to space mission requirements and enabling both reconfiguration and failure mitigation. For example, the authors employed the on-board processor to create a vision-based navigation system. The system has the ability to autonomously adapt to different stages of a mission or unexpected situations, highlighting the practical benefits of reconfigurable MPSoCs for real-world space missions. In addition, the research presented in [83] indicates that the use of Digital Signal Processors (DSPs) greatly enhances the ability to handle errors while performing crucial calculations, in contrast to the use of Look-Up Tables (LUTs). This results in a significant decrease in downtime, with reductions of 95% and 65% for certain activities, thus reinforcing the argument for using reconfigurable MPSoCs in space applications.
Cost-effective solutions are frequently necessary for new space missions. The authors of [84] aim to enhance the efficacy of star trackers in satellites, which is crucial for preserving orientation. They accomplish this by employing reconfigurable MPSoC to accelerate image processing algorithms, which is a more cost-effective method than conventional methods. Their accomplishments were remarkable: they were able to accelerate computations by 95% and process 4-megapixel images at a rate of over 24 frames per second through the use of ingenious design techniques. This optimization ensures that the system operates in real-time and remains accurate, rendering it a viable option for space programs that prioritize cost.
Figure 4: The proposed EO on-board processing for semantic communication inspired by [82].
V-B3 Runtime and modular reconfiguration on-board processing
The avionics and space industries are experiencing an increase in the popularity of on-board processors. This technology, which was first implemented on the FedSat mission [85], is currently employed to resolve errors and modify functions in mid-flight, as seen in Cibola. This dual-purpose reconfiguration is advantageous for electronics that operate in challenging environments. This research establishes a spacecraft system that is adaptable by utilizing the reconfigurable ARTICo3 architecture [86]. ARTICo3 enables the hardware to be reprogrammed during flight to ensure fault tolerance and adaptable performance. The system seamlessly integrates with a space-ready operating system (RTEMS) and includes error-handling capabilities. Additionally, it facilitates both mission-specific duties and general tasks (e.g., defect management and reconfiguration). For instance, the researchers of [86] demonstrated the system’s adaptability by incorporating vision processing that adjusts to various mission phases. This method illustrates a promising avenue for the development of spaceborne systems that are both flexible and dependable. It is worth noting that ESA’s study presented in [87] expands the limits of on-board partial reconfiguration using the BRAVE platform developed by NanoExplore [88]. BRAVE FPGAs are purpose-built for space applications, with a radiation-resistant material that ensures their reliability in challenging situations. An important advancement is the use of a SpaceWire interface for reconfiguration, which is a specialized space communication standard that may provide benefits compared to conventional JTAG approaches as investigated in an independent study [89]. This exemplifies the continuous progress in reconfigurable on-board processing for versatile and reliable spacecraft.
As illustrated in Fig.4, we propose implementing an on-board processing for semantic communication EO based on the research the study presented in [82] using a Zynq UltraScale+ MPSoC. The system integrates versatile Application Processing Unit (APU) to handle a wide range of activities, dedicated Real-Time Processing Unit (RPU) for time-sensitive operations, and customizable reconfigurable Programmable Logic (PL). Real-Time Executive for Multiprocessor Systems (RTEMS), a real-time operating system, effectively controls all tasks and guarantees prompt execution including semantic task schedule. The RPU and APU operate the essential application of the satellite, such as navigation, in addition to performing support activities including managing errors, controlling reconfigurations, and communicating with ground base station. This platform focuses on enhancing the fault tolerance of spacecraft computers via the use of a specialized Cortex-R5 processor-based RPU. The RPU utilizes lock-step execution, which involves completing computations twice in order to identify mistakes. The attainment of reconfigurability is made possible by the use of the Zynq UltraScale+ MPSoC hardware component, which enables adjustments to be made during missions. The Xilinx Soft Error Mitigation (SEM) IP proactively detects and rectifies problems induced by radiation. The RPU functions as the central hub, overseeing both reconfiguration and error correction. The fault-tolerant design incorporates redundant processing, modular hardware, and on-chip correction to enable dependable and adaptive spaceborne computing. Efficient communication is achieved via the use of separate AXI4 busses. The use of distinct AXI4 busses enables effective communication within this resilient design. The integration of redundant processing, adaptable hardware featuring deep learning accelerator capabilities, and on-chip error correction enables the development of resilient and flexible computing platforms for spacecraft. This platform is essential for ensuring the reliable execution of deep learning or other resource-intensive tasks in challenging space environments.
V-C Numerical Results for on-board EO deep learning
Digital Task-oriented Joint Source-Channel Coding (DT-JSCC) [14] enables the seamless integration of digital communication and DJSCC. The initial modulation, known as uniform modulation, operates like an adaptable spring, adjusting the spacing between data points (constellations) during transmission to better align with the specific characteristics of the transmitted image data. This is the most suitable option for semantic communication in the aforementioned circumstance.
V-C1 Physical layer channel model
The signals that were transmitted and received are denoted by $X$ and $Y$, respectively. The general form of the channel equation is $Y=H\cdot X+N$, where $H$ is the channel gain matrix and $N$ is the noise term. Below are specific formulae for different sorts of channels. $Y=X+N_{\text{AWGN}}$, where $N_{\text{AWGN}}\sim\mathcal{N}(0,\sigma^{2})$, because $H=1$ due to the absence of fading and has a Gaussian noise variance of $\sigma^{2}$. However, Rician distribution characterizes the channel gain $H$. The expression $Y=H_{\text{Rician}}\cdot X+N_{\text{AWGN}}$ is Rician-distributed, and the Rician $K_{r}$-factor defines $H_{\text{Rician}}$.
V-C2 Simulation results using EuroSAT dataset
Top-1 accuracy utilizing DT-JSCC for EuroSAT data under various channel circumstances is shown in Fig. 5. The performance of the 16PSK/16APSK modulation technique at various PSNR (Peak Signal-to-Noise Ratio) levels is shown in Fig. 5. With varying K-factor values (32, 64, 128), we demonstrate accuracy for the Rician and AWGN (Additive White Gaussian Noise) channels; in general, a larger K-factor indicates better channel quality. Additionally, for these K values, we compare accuracy in the Rician channel. The figures show that accuracy increases with rising PSNR in all conditions, that AWGN channels outperform Rician in general, and that greater K-factors result in more steady performance.
@@ V-C2 Simulation results using EuroSAT dataset
In comparing the confusion matrices of Fig.6, we can observe some key performance differences over the AWGN channel where PSNR=14dB. For k=128, the classification accuracy across most categories is notably high, with certain categories like ”Forest” showing a strong prediction rate of 99.21%, while there are minor misclassifications in categories like ”PermanentCrop” and ”Pasture” with accuracies of 93.71% and 94.59% respectively. Meanwhile, the case of K=32 also performs well but demonstrates slightly different classification behaviors, such as a stronger prediction for ”River” at 98.05% but more substantial confusion in categories like ”Forest” and ”Residential,” where there are larger misclassification rates with accuracies of 95.44% and 90.88% compared to the K=128 performance. The identical trend of K=32 and K=128 can be observed over the 16APSK Rician channel.
@@
As illustrated in the confusion matrices of Fig.7, we can observe some key performance differences for DT-JSCC using K=128, the classification accuracy across most categories is notably high, with certain categories like ”Forest” showing a strong prediction rate of 99.88%, while there are minor misclassifications in categories like ”PermanentCrop” and ”Pasture” with accuracies of 98.34%.
@@ VI Conclusion
In this paper, we explored the critical role of semantic inference models and deep learning in enhancing EO systems, focusing on the integration of data-driven modeling approaches. With the rise of big EO data and the increasing complexity of satellite imagery, conventional methods struggle to meet the demands of modern applications such as environmental monitoring, disaster management, and resource allocation. By utilizing advanced deep learning techniques, we have illustrated the potential to overcome key challenges, including the scarcity of labeled data, the variability in data sources, and the intricacies of multi-dimensional datasets. The incorporation of semantic inference enables systems to extract meaningful insights from vast and complex datasets, paving the way for more accurate decision-making in both civilian and critical operations. The advancements in on-board processing hardware demonstrate the increasing feasibility of efficient and adaptable computation for space missions. The evolution of SRAM-based FPGAs, radiation-hardened processors, and reconfigurable MPSoCs has opened new avenues for enhancing spacecraft resilience, adaptability, and task performance in real time. These systems ensure precise handling of computational tasks while mitigating the effects of space radiation, a critical capability for sustained mission success. Ultimately, on-board processing solutions promise improved data handling and communication efficiency, aligning with the goals of modern satellite technology and paving the way for sophisticated applications in space exploration.